Part 2. Cross rates

The highest bid and lowest ask in the NY currency market for the listed currencies:

Australian dollar

AUD/USD

0.7190

Chinese yuan

USD/CNY

6.7473

Danish krone

USD/DKK

6.5451

Euro

EUR/USD

1.1415

Jamaican dollar

USD/JMD

132.88

Japanese yen

USD/JPY

109.55

New Zealand dollar

NZD/USD

0.6839

West African franc

USD/XOF

279.50

1 a. What is the implied cross rate for exchanging Australian dollars and New Zealand dollars? What is the base currency in this cross rate?

b. If a customer sells A$1m in exchange for NZ$, how many NZ$ will the customer receive?

2 a. What is the implied cross rate for exchanging Chinese yuan and Danish krone? Specify which currency is the base currency in your answer.

b. If a customer sells Kr 1m in exchange for ¥, how many ¥ will the customer receive?
Answer #1

1a) To Calculate Cross Rate for exchanging Australian Dollar and New Zealand Dollar we need

AUD/USD = 0.7190

NZD/USD = 0.6839

AUD/NZD = AUD/USD * 1/NZD/USD

= 0.7190 * 1/0.6839

= 0.7190 * 1.4622

= 1.0513

or AUD/NZD = 1.0513 where base currency is New Zealand Dollar

1b) If a Customer sells AUD 1 Million he will get

= 1000000 * 1.0513

= 1051300 NZD

2a) To Calculate Cross Rate for exchanging Chinese Yuan and Danish Krone we need

USD/CNY = 6.7473

USD/DKK = 6.5451

CNY/DKK = 1/USD/CNY * USD/DKK

= 1/6.7473 * 6.5451

= 0.1482 * 6.5451

= 0.9700

or CNY/DKK = 0.9700 where base currency is Danish Krone

1b) If a Customer sells DKK 1 Million he will get

= 1000000 / 0.9700

= 1030928 CNY